Given a possible triangle ABC with a=16, b=25, and A= 22(degrees), find a possible value for B. Round to the nearest tenth.
Answer by stanbon(75887)

Given a possible triangle ABC with a=16, b=25, and A= 22(degrees), find a possible value for B. Round to the nearest tenth.
----
sin(B)/25 = sin(22)/16
---
sin(B) = (25/16)sin(22)
---
sin(B) = 0.5853
B = sin^-1(0.5853) = 35.8 degrees
====================================
Cheers,
Stan H.
==========